OpenStudy (anonymous):

how do i find all the minimum values of x+sin^2 x with the domain of pi/6 to 5pi/6?

OpenStudy (perl):

first find the critical points solve f ' (x) = 0

OpenStudy (anonymous):

i found that the derivative is 1+ 2sincos and that the critical point is 3pi/4

OpenStudy (perl):

ok

OpenStudy (anonymous):

now what?

OpenStudy (perl):

now you test the critical points and endpoints

OpenStudy (anonymous):

i plug them back into the original equation?

OpenStudy (perl):

the critical point is not 3pi/4

OpenStudy (perl):

1 + 2sinx cos x = 0 1 + sin(2x) = 0 sin(2x) = -1 2x= arcsin ( -1) x = arcsin(-1) / 2 #######don't forget to divide by 2 #######

OpenStudy (perl):

x = ( 3pi/4 + 2pi*n ) / 2 x = 3pi/8 + pi*n , and since we want domain [pi/6, 5pi/6] x = 3pi/8 is our critical point

OpenStudy (perl):

ok so far? did that make sense

OpenStudy (anonymous):

yes

OpenStudy (perl):

ok now we use the Absolute maxima theorem If f(c) is an absolute maximum/minimum, then c is either a critical point or an endpoint.

OpenStudy (perl):

so evaluate f(pi/6)= f(3pi/8)= f(5pi/6) =

OpenStudy (anonymous):

and is the lowest of those the minimum?

OpenStudy (perl):

correct

OpenStudy (anonymous):

thanks!